OpenStudy (anonymous):

c=6x-4y If I were to plug in points.. would it look something like this?

OpenStudy (anonymous):

c = 6(0,0)(5,0) - 4(0,3) ? Here's the link for the intersection points for the following question. https://www.desmos.com/calculator/ygtpnkixcp

OpenStudy (anonymous):

Given the system of constraints, name all vertices of the feasible region. Then find the maximum value of the given objective function. x>=0 y>=0 y<=1/3x+3 5>=y+x Objective Function: C = 6x – 4y